The Number of Condorcet Scenarios with 2n (resp. 2n+1) Voters and Three Candidates is 2\binom{n+2}{5} (resp. 2\binom{n+4}{5})

Shalosh B. EKHAD^{1}

1 c/o D. Zeilberger, Department of Mathematics, Rutgers University (New Brunswick), Hill Center-Busch Campus, 110 Frelinghuysen Rd., Piscataway, NJ 08854-8019, USA. c/o zeilberg at math dot rutgers dot edu ,

http://www.math.rutgers.edu/~zeilberg/ . June 4, 2012.